# Text-Queried Audio Source Separation via Hierarchical Modeling

### v1

@@ III-B1 Local semantic representation
We leverage the AudioMAE [45], a self-supervised learned audio representation, for the local semantic feature $S$. Unlike discrete neural codec codes that prioritizes acoustic fidelity but lacks semantic representativeness, or contrastively learned cross-modal features CLAP that lose details, AudioMAE balances acoustic details and semantic information by predicting masked acoustic features. This makes it a good local-semantic representation, clustering semantically similar audio events while retaining spatial-temporal structure.
AudioMAE takes a Vision Transformer (ViT)-based encoder-decoder structure. It takes the log mel spectrogram $X\in\mathbb{R}^{T\times F}$ of an audio signal $x$ as input, splits it into $P\times P$ patches, and embeds them into a $C_{s}$-dimensional latent space, yielding features of a shape $\frac{T}{P}\times\frac{F}{P}\times C_{s}$. $T$ and $F$ are the number of frames and mel-frequency bins, respectively. Its masked autoencoder design uses an asymmetric structure, pairing large encoders with small decoders and scales well for linear probing, thanks to a high masking ratio that reduces encoding complexity [44]. Following MW-MAE [48], we adopt a small 4-layer transformer-based decoder with a width of 384 and 8 attention heads and pretrain it on AudioSet [23].
@@ III-B2 Global semantic representation
While AudioMAE captures both semantic and acoustic details, it lacks alignment with text feature space, limiting cross-modal understanding. To address this, we introduce Q-Audio as the cross-modal global semantic feature extractor, inspired by BLIP-2 [46]. As depicted in Figure 2 (a), Q-Audio bridges features from a frozen local semantic audio encoder (our pretrained AudioMAE) and a frozen FLAN-T5 text encoder [49] into a shared space. It comprises two feature extractors that utilize a transformer structure and share self-attention layers. The audio part extracts semantic components from AudioMAE via a self-attention followed by a cross-attention, and the text feature extractor utilizes only self-attention layers. We adopt one learnable query embedding to interact with the frozen audio features through cross-attention layers. The query also interacts with the text branch in self-attention layers through masking. This structure distills global semantic information from the local-semantic representation while alleviating text comprehension demands through FLAN-T5, which is different from the original BLIP-2 design [46]. This design makes the extractor an audio-language aligner with a much lower complexity than the original one. What’s more, extracting global feature $G\in\mathbb{R}^{1\times C_{g}}$ from the local-semantic representation $S\in\mathbb{R}^{\frac{T}{P}\times\frac{F}{P}\times C_{s}}$ further ensures the correlation and consistency between two-level representations, better facilitating our hierarchical modeling and separation design.
During the learning phase, the main objective is to enable the learnable query to extract audio representations that are most informative for the corresponding text input. We adopt the original optimization design in [46] with three training objectives: audio-language contrastive learning, audio-language matching, and audio-grounded text generation, each leveraging a different attention masking mechanism to regulate the interaction between the learnable query and the text branch in self-attention layers. As shown in Figure 2, we use the audio branch of Q-Audio as the global-semantic audio encoder, and the text branch as the text encoder for the two-level separation.
@@ III-C Global-semantic separation
The first stage $\mathcal{H}_{1}:(x_{mix},\mathcal{T})\mapsto\hat{G}$ of our separation framework operates in the Q-Audio feature space to separate semantic representations of the target audio. We use Q-Audio as the mixture audio and text encoders and employ a 6-layer non-autoregressive (NAR) transformer to predict the target semantic feature $G_{tgt}\in\mathbb{R}^{1\times C_{g}}$. This prediction is conditioned on the concatenation of the mixture audio feature $G_{mix}\in\mathbb{R}^{1\times C_{g}}$, the text feature $T_{cap}\in\mathbb{R}^{N\times C_{t}}$, and the task token $T_{task}$. $N$ is the sequence length of text tokens and $C_{t}$ is the token dimension. The semantic alignment between audio and text feature inputs enables the transformer to accurately model audio-text relationships and guide the separation process.
@@ III-D Local-semantic separation
The second stage $\mathcal{H}_{2}:(x_{mix},\mathcal{T},\hat{G})\mapsto\hat{S}$ operates in the more fine-grained AudioMAE feature space. This stage takes the mixture audio feature $S_{mix}\in\mathbb{R}^{\frac{T}{P}\times\frac{F}{P}\times C_{s}}$ by AudioMAE encoder as input to predict the target feature $S_{tgt}\in\mathbb{R}^{\frac{T}{P}\times\frac{F}{P}\times C_{s}}$. The previous-stage output $\hat{G}$, text feature $T_{cap}\in\mathbb{R}^{N\times C_{t}}$, and the task token $T_{task}$ are concatenated as the conditioning input, as shown in Figure 2. During training, ground-truth global audio feature $G_{tgt}$ is used instead of the prediction $\hat{G}$.
@@
$\mathcal{L}_{local}=||\hat{S}-S_{tgt}||_{1},$ (8)
Where $\hat{S}$ is the predicted output. To mitigate possible error propagation between two stages, we further conduct joint fine-tuning by Equations 7 and 8 after optimizing each stage independently. We introduce a switcher to control the conditioning input of the second stage, selecting ground-truth $G_{tgt}$ with a probability of $P_{gt}$, and prediction $\hat{G}$ with a probability of $P_{pred}=1-P_{gt}$. This joint fine-tuning optimizes the two stages end-to-end, minimizing inconsistencies and improving separation accuracy.
III-E Acoustic decoder
After the two-stage separation, the last part $\mathcal{H}_{3}:\hat{S}\mapsto x_{tgt}$, the acoustic decoder, aims to reconstruct the waveform from the predicted local semantic feature. In light of the power of generative models, we leverage auto-regressive transformer for this stage. Specifically, we encode $x_{tgt}$ into discrete acoustic tokens $A$ using a neural audio codec, and leverage an autoregressive transformer to generate audio tokens conditioned on $\hat{S}$, followed by a neural codec decoder to reconstruct the waveform, similar to that in UniCodec [50].
@@ III-E2 Reconstruction via TF-Codec
Figure 3: Overview of the acoustic decoder. The autoregressive transformer generates acoustic tokens by our neural codec TF-Codec, conditioned on local semantic features.
@@
TF-Codec [51] is a low-latency neural speech codec designed for high-quality speech at low bitrates. We retrain a non-predictive version of TF-Codec to adapt to general audio for acoustic tokens. It employs a VQ-VAE framework, including an encoder, a group vector quantizer, and a decoder optimized end to end. To ensure high perceptual quality for diverse general audio, we apply adversarial training with a multi-scale mel discriminator [19] to replace the original single-scale frequency-domain discriminator in [51]. Other training losses are the same as that in [51]. Instead of training from scratch, we finetune from the pre-trained TF-Codec speech codec model for better performance. Similar to UniCodec [50], in autoregressive token generation of Equation 9, all groups of $t$-th step are simultaneously generated in a single stage, leading to a short token length.
III-F Instruction semantic parsing
In text-guided audio separation, users typically provide descriptive prompts to retain some audio components or remove some undesired audio events. This instruction is arbitrary and may target at single or multiple sound sources with specified characteristics. To facilitate this open-vocabulary text prompts with bidirectional operations, our framework introduces dual separation operations: removal and extraction, allowing the model to isolate target audio through complementary approaches. We also utilize pre-trained LLMs to parse text prompts into this task type and a target audio description. For example, an input “Could you help me separate the sound of a dog barking and the background music?” is decomposed into a task type “extraction” and an audio description “a dog barking and the background music”, enabling distinct processing of operation intent and acoustic context.
@@ III-F Instruction semantic parsing
This process offers two critical advantages. It enhances the prompt understanding capability of the separation model. What’s more, the decomposition of event description from operation types makes it feasible to leverage audio-language contrastive learning to extract aligned audio-text features, improving the separation model’s ability to capture cross-modal correlations. We introduce a dual-channel text conditioning mechanism for this decomposed processing. The task type is encoded using pretrained token representations from T5’s vocabulary, producing continuous task tokens $T_{task}\in\mathbb{R}^{L\times C_{t}}$, where $L$ is the token length and $C_{t}$ is the embedding dimension. Simultaneously, the event description is processed by the semantic text encoder, denoted as $T_{cap}\in\mathbb{R}^{N\times C_{t}}$, where $N$ is the length of the caption sequence. These two streams of tokens are concatenated along the sequence dimension to form the final text conditioning input $[T_{task},T_{cap}]\in\mathbb{R}^{(L+N)\times C_{t}}$, as shown in Figure 2 (b).
@@ IV-A1 Semantic separation
We utilize audio data from AudioCaps [52], ESC-50 [53], Clotho v2 [54], FSD50K [55] and WavCaps [56] due to their relatively accurate and diverse labels or captions. Each audio contains single or multiple audio events with or without overlaps. AudioCaps is a subset of AudioSet [23] with handcrafted captions and we use 45K audio clips. Clotho v2 provides human-annotated captions, and we use 3,839 training clips, each accompanied by five captions. ESC-50 comprises 2,000 environmental audio recordings, evenly distributed across 50 semantic classes, including natural sounds, non-speech human sounds, domestic sounds, and urban noises. FSD50K contains approximately 40K audio clips with 200 classes, varying in duration from 0.3 to 30 seconds. WavCaps includes four subsets: AudioSet_SL, Freesound, SoundBible, and BBC Sound Effect. We use the entire AudioSet_SL subset, exclude Freesound with inaccurate captions, and filter the SoundBible and BBC Sound Effect subsets by removing audios longer than 60 seconds. Ultimately, we select 121K audio clips from WavCaps. All audios are resampled to 16 kHz and we normalize all of them to 10 seconds by cropping or padding. In total, we compile 230K ten-second audio clips with class labels or captions as the source data for generating mixtures.
@@
For semantic separation, we employ triplet data (mixed audio, target audio, instruction) for supervised training. Instructions are generated using two task types and audio descriptions. We generate bidirectional instructions for each triplet and randomly select one during training, balancing the model’s ability to interpret both target and undesired audio events. Given that most source audios contain multiple audio events, when randomly selecting two samples to create each mixture, we follow the rule that the separated two audios do not have any overlap in audio classes. As text-audio pairs from various datasets feature diverse audio descriptions (e.g. FSD50K and ESC-50 provide different fixed sets of class labels, while AudioCaps, WavCaps and Clotho v2 offer detailed captions), it is necessary to unify the class labels to distinguish audio event types during audio mixing. We employ GPT-4111https://chatgpt.com/ [57] to semantically categorize different audio types of different datasets and summarize audio event tags from captions. For example, “rain” and “raindrop” are the same type and GPT-4 extracts tags “man speech, thump, frog croak” from the original caption “An adult male speaks while thumps occur in the background, then frogs croak.” It is noted that these unified tags are only used in audio mixing to avoid mixing audios including the same categories. The instruction part will keep the original caption and labels to preserve the diversity. In total, we create 1.7M mixture-target audio pairs for training, containing approximately 4500 hours.
@@ IV-A2 Q-Audio and AudioMAE
To pretrain the Q-Audio module, we use the same datasets as that in Section IV-A1, i.e. AudioCaps, ESC-50, Clotho v2, FSD50K and WavCaps, which provide detailed captions or class labels paired with each audio. Following the approaches of AudioMAE and MW-MAE [48], we train the MAE model using AudioSet-2M [23], encompassing both balanced and unbalanced subsets.
IV-A3 Acoustic decoder
The acoustic decoder is trained using audios only without text labels. We gather data from AudioSet for general sounds, the Jamendo dataset for music [58], and a multilingual speech dataset with 6k hours for clean speech [50] that combines the Libri-Light medium subset [59], the multilingual speech subsets of the ICASSP 2021 DNS Challenge corpus [60], and AVSpeech [61]. All audios are resampled to 16 kHz. We randomly sample 2,000 hours from each category. During TF-Codec training, audios are cropped to 3-second segments. For autoregressive audio token generation module, we set the maximum duration of each training audio to 10 seconds and perform padding masks on audio shorter than 10 seconds.
@@ IV-B1 Evaluation benchmark
We compile evaluation data from the test sets of AudioCaps, FSD50K, and Clotho v2, for general audio assessment. The mixing strategy is the same as that used for training. As shown in Table I, we also create a “3 Sets” that combine three datasets with an equal number of randomly sampled audio clips from each source dataset. Mixing sources from different datasets can account for distribution biases, leading to better approximations of real-world recordings. The average SNR between two components of the mixture is controlled within [-15dB, 15dB].

### v2

@@ III-B1 Local semantic representation
We leverage the AudioMAE [huang2022masked], a self-supervised learned audio representation, for the local semantic feature $S$. Unlike discrete neural codec codes that prioritizes acoustic fidelity but lacks semantic representativeness [borsos2023audiolm, jiang2024universal], or contrastively learned cross-modal features CLAP that lose details, AudioMAE balances acoustic details and semantic information by predicting masked acoustic features. This makes it a good local-semantic representation, clustering semantically similar audio events while retaining spatial-temporal structure, as also discussed in AudioLDM 2 [liu2024audioldm].
AudioMAE takes a Vision Transformer (ViT)-based encoder-decoder structure. It takes the log mel spectrogram $X\in\mathbb{R}^{T\times F}$ of an audio signal $x$ as input, splits it into $P\times P$ patches, and embeds them into a $C_{s}$-dimensional latent space, yielding features of a shape $\frac{T}{P}\times\frac{F}{P}\times C_{s}$. $T$ and $F$ are the number of frames and mel-frequency bins, respectively. Its masked autoencoder design uses an asymmetric structure, pairing large encoders with small decoders and scales well for linear probing, thanks to a high masking ratio that reduces encoding complexity [he2022masked]. Following MW-MAE [yadav2023masked], we adopt a small 4-layer transformer-based decoder with a width of 384 and 8 attention heads and pretrain it on AudioSet [gemmeke2017audio].
@@ III-B2 Global semantic representation
While AudioMAE captures both semantic and acoustic details, it lacks alignment with text feature space, limiting cross-modal understanding. To address this, we introduce Q-Audio as the cross-modal global semantic feature extractor, inspired by BLIP-2 [li2023blip]. As depicted in Figure 2 (a), Q-Audio bridges features from a frozen local semantic audio encoder (our pretrained AudioMAE) and a frozen FLAN-T5 text encoder [chung2024scaling] into a shared space. It comprises two feature extractors that utilize a transformer structure and share self-attention layers. The audio part extracts semantic components from AudioMAE via a self-attention followed by a cross-attention, and the text feature extractor utilizes only self-attention layers. We adopt one learnable query embedding, a trainable latent vector that selectively aggregates and distills information from input representations [li2023blip]. This embedding interacts with the frozen audio features via cross-attention mechanisms, thereby facilitating efficient extraction of global semantic cues aligned with the textual query. The query also interacts with the text branch in self-attention layers through masking. This structure extracts global semantic information from the local-semantic representation while alleviating text comprehension demands through FLAN-T5, which is different from the original BLIP-2 design [li2023blip]. This design makes the extractor an audio-language aligner with a much lower complexity than the original one. What’s more, extracting global feature $G\in\mathbb{R}^{1\times C_{g}}$ from the local-semantic representation $S\in\mathbb{R}^{\frac{T}{P}\times\frac{F}{P}\times C_{s}}$ further ensures the correlation and consistency between two-level representations, better facilitating our hierarchical modeling and separation design.
During the learning phase, the main objective is to enable the learnable query to extract audio representations that are most informative for the corresponding text input. We adopt the original optimization design in [li2023blip] with three training objectives: audio-language contrastive learning, audio-language matching, and audio-grounded text generation, each leveraging a different attention masking mechanism to regulate the interaction between the learnable query and the text branch in self-attention layers. As shown in Figure 2, we use the audio branch of Q-Audio as the global-semantic audio encoder, and the text branch as the text encoder for the two-level separation.
@@ III-C Global-semantic separation
The first stage $\mathcal{H}_{1}:(x_{mix},\mathcal{T})\mapsto\hat{G}$ of our separation framework operates in the Q-Audio feature space to separate semantic representations of the target audio. We use Q-Audio as the mixture audio and text encoders and employ a 6-layer non-autoregressive (NAR) transformer to predict the target semantic feature $G_{tgt}\in\mathbb{R}^{1\times C_{g}}$. Here, $G_{tgt}$ denotes the global semantic feature extracted by Q-Audio taking the ground-truth target signal $x_{tgt}$ as input. This prediction is conditioned on the concatenation of the mixture audio feature $G_{mix}\in\mathbb{R}^{1\times C_{g}}$, the text feature $T_{cap}\in\mathbb{R}^{N\times C_{t}}$, and the task token $T_{task}$. $N$ is the sequence length of text tokens and $C_{t}$ is the token dimension. The semantic alignment between audio and text feature inputs enables the transformer to accurately model audio-text relationships and guide the separation process.
@@ III-D Local-semantic separation
The second stage $\mathcal{H}_{2}:(x_{mix},\mathcal{T},\hat{G})\mapsto\hat{S}$ operates in the more fine-grained AudioMAE feature space. This stage takes the mixture audio feature $S_{mix}\in\mathbb{R}^{\frac{T}{P}\times\frac{F}{P}\times C_{s}}$ by AudioMAE encoder as input to predict the target AudioMAE feature $S_{tgt}\in\mathbb{R}^{\frac{T}{P}\times\frac{F}{P}\times C_{s}}$ of the target audio $x_{tgt}$. The previous-stage output $\hat{G}$, text feature $T_{cap}\in\mathbb{R}^{N\times C_{t}}$, and the task token $T_{task}$ are concatenated as the conditioning input, as shown in Figure 2. During training, ground-truth global audio feature $G_{tgt}$ is used instead of the prediction $\hat{G}$.
@@
[…]
Where $\hat{S}$ is the predicted output. To mitigate possible error propagation between two stages, we further conduct joint fine-tuning by Equations 7 and 8 after optimizing each stage independently. We introduce a switcher to control the conditioning input of the second stage, selecting ground-truth $G_{tgt}$ with a probability of $P_{gt}$, and prediction $\hat{G}$ with a probability of $P_{pred}=1-P_{gt}$, similar to that in AudioLDM 2 [liu2024audioldm]. This joint fine-tuning optimizes the two stages end-to-end, minimizing inconsistencies and improving separation accuracy.
III-E Semantic-to-Acoustic and Acoustic decoder
After the two-stage separation, the last part $\mathcal{H}_{3}:\hat{S}\mapsto x_{tgt}$, the decoder, aims to reconstruct the waveform from the predicted local semantic feature. In light of the power of generative models, we leverage auto-regressive transformer for this stage. Specifically, we encode $x_{tgt}$ into discrete acoustic tokens $A$ using a neural audio codec, and leverage an autoregressive transformer to generate audio tokens conditioned on $\hat{S}$, followed by a neural codec decoder to reconstruct the waveform, similar to that in UniCodec [jiang2024universal].
@@ III-E2 Reconstruction via TF-Codec
Figure 3: Overview of the semantic-to-acoustic and acoustic decoder. The autoregressive transformer generates acoustic tokens by our neural codec TF-Codec, conditioned on local semantic features.
@@
TF-Codec [jiang2023latent] is a low-latency neural speech codec designed for high-quality speech at low bitrates. We retrain a non-predictive version of TF-Codec to adapt to general audio for acoustic tokens. It employs a VQ-VAE framework, including an encoder, a group vector quantizer, and a decoder optimized end to end. To ensure high perceptual quality for diverse general audio, we apply adversarial training with a multi-scale mel discriminator [yang2023uniaudio] to replace the original single-scale frequency-domain discriminator in [jiang2023latent]. Other training losses are the same as that in [jiang2023latent]. Instead of training from scratch, we finetune from the pre-trained TF-Codec speech codec model for better performance. Similar to UniCodec [jiang2024universal], in autoregressive token generation of Equation 9, all groups of $t$-th step are simultaneously generated in a single stage, leading to a short token length.
III-F Bidirectional Query Instruction
In text-guided audio separation, users typically provide descriptive prompts to retain certain audio components or remove undesired audio events. To support bidirectional operations, our framework formulates the separation task into two complementary modes: removal and extraction. Each instruction is decomposed into a task type, indicating the operation, and an audio event description, specifying the relevant sound sources.
@@ III-F Bidirectional Query Instruction
We introduce a dual-channel text conditioning mechanism for this decomposed processing. The task type is encoded using pretrained token representations from T5’s vocabulary, producing continuous task tokens $T_{task}\in\mathbb{R}^{L\times C_{t}}$, where $L$ is the token length and $C_{t}$ is the embedding dimension. Simultaneously, the event description is processed by the semantic text encoder, denoted as $T_{cap}\in\mathbb{R}^{N\times C_{t}}$, where $N$ is the length of the caption sequence. These two streams of tokens are concatenated along the sequence dimension to form the final text conditioning input $T_{task},T_{cap}]\in\mathbb{R}^{(L+N)\times C_{t}}$, as shown in Figure 2 (b).
This design offers two critical advantages. First, it enhances the query understanding capability and extensibility of the separation model by structurally disentangling the operation type from the audio event description. Second, this decomposition enables the use of audio-language contrastive learning to extract aligned audio-text features, improving the model’s ability to capture cross-modal correlations. In practical cases, for arbitrary instructions (e.g. Can you help me separate the sound of a dog barking and the background music?), we can utilize large language models (LLM) to assist in parsing natural language prompts into the structured format by our framework, as shown in Figure 4.
@@ IV-A1 Semantic separation
We utilize audio data from AudioCaps [kim2019audiocaps], ESC-50 [piczak2015esc], Clotho v2 [drossos2020clotho], FSD50K [fonseca2021fsd50k] and WavCaps [mei2024wavcaps] due to their relatively accurate and diverse labels or captions. Each audio contains single or multiple audio events with or without overlaps. AudioCaps is a subset of AudioSet [gemmeke2017audio] with handcrafted captions and we use 45K audio clips. Clotho v2 provides human-annotated captions, and we use 3,839 training clips, each accompanied by five captions. ESC-50 comprises 2,000 environmental audio recordings, evenly distributed across 50 semantic classes, including natural sounds, non-speech human sounds, domestic sounds, and urban noises. FSD50K contains approximately 40K audio clips with 200 classes, varying in duration from 0.3 to 30 seconds. WavCaps includes four subsets: AudioSet_SL, Freesound, SoundBible, and BBC Sound Effect. We use the entire AudioSet_SL subset, exclude Freesound with inaccurate captions, and filter the SoundBible and BBC Sound Effect subsets by removing audios longer than 60 seconds. Ultimately, we select 121K audio clips from WavCaps. All audios are resampled to 16 kHz and we normalize all of them to 10 seconds by cropping or padding. In total, we compile 230K ten-second audio clips with class labels or captions as the source data for generating mixtures.
@@
For semantic separation, we employ triplet data (mixed audio, target audio, instruction) for supervised training. Instructions are generated using two task types and audio descriptions. We generate bidirectional instructions for each triplet and randomly select one during training, balancing the model’s ability to interpret both target and undesired audio events. For example, for a mixture with two sources (A and B), we generate two types of instructions: (1) an extraction instruction targeting one source (e.g., “extract A”), and (2) a removal instruction targeting the other source (e.g., “remove B”). Given that most source audios contain multiple audio events, when randomly selecting two samples to create each mixture, we follow the rule that the separated two audios do not have any overlap in audio classes. As text-audio pairs from various datasets feature diverse audio descriptions (e.g. FSD50K and ESC-50 provide different fixed sets of class labels, while AudioCaps, WavCaps and Clotho v2 offer detailed captions), it is necessary to unify the class labels to distinguish audio event types during audio mixing. We employ GPT-4111https://chatgpt.com/ [achiam2023gpt] to semantically categorize different audio types of different datasets and summarize audio event tags from captions. For example, “rain” and “raindrop” are the same type and GPT-4 extracts tags “man speech, thump, frog croak” from the original caption “An adult male speaks while thumps occur in the background, then frogs croak.” It is noted that these unified tags are only used in audio mixing to avoid mixing audios including the same categories. The instruction part will keep the original caption and labels to preserve the diversity. In total, we create 1.7M mixture-target audio pairs for training, containing approximately 4500 hours.
@@ IV-A2 Q-Audio and AudioMAE
To pretrain the Q-Audio module, we use the same datasets as that in Section IV-A1, i.e. AudioCaps, ESC-50, Clotho v2, FSD50K and WavCaps, which provide detailed captions or class labels paired with each audio. Following the approaches of AudioMAE and MW-MAE [yadav2023masked], we train the MAE model using AudioSet-2M [gemmeke2017audio], encompassing both balanced and unbalanced subsets.
IV-A3 Semantic-to-acoustic and acoustic decoder
The decoder is trained using audios only without text labels. We gather data from AudioSet for general sounds, the Jamendo dataset for music [bogdanov2019mtg], and a multilingual speech dataset with 6k hours for clean speech [jiang2024universal] that combines the Libri-Light medium subset [kahn2020libri], the multilingual speech subsets of the ICASSP 2021 DNS Challenge corpus [reddy2021icassp], and AVSpeech [ephrat2018looking]. All audios are resampled to 16 kHz. We randomly sample 2,000 hours from each category. During TF-Codec training, audios are cropped to 3-second segments. For autoregressive audio token generation module, we set the maximum duration of each training audio to 10 seconds and perform padding masks on audio shorter than 10 seconds.
@@ IV-B1 Evaluation benchmark
We compile evaluation data from the test sets of AudioCaps, FSD50K, and Clotho v2, for general audio assessment. The mixing strategy is the same as that used for training. As shown in Table I, we also create a “3 Sets” that combine three datasets with an equal number of randomly sampled audio clips from each source dataset. Mixing sources from different datasets can account for distribution biases, leading to better approximations of real-world recordings. The average SNR between two components of the mixture is controlled within [-15dB, 15dB]. It is noted that we utilize the simple “extract ”and “removal” instruction templates with original captions of the source datasets to ensure a fair comparison.